Integration by Partial Fractions


1.) ∫dz/z+z^3


2.) ∫(2s+1)ds/s^2(s^2+1)



1
Expert's answer
2021-05-07T14:37:28-0400

1.) "\int \dfrac{dz}{z+z^3}"


Partial fraction can be written as,


"f(z) = \dfrac{1}{z} - \dfrac{z}{z^2+1}"


"f(z) = \dfrac{1}{z} - \dfrac{1}{2} \dfrac{1}{(z+i)}+ \dfrac{1}{2} \dfrac{1}{(z-i)}"


Integrating both sides,


"\int f(z)dz = logz - \dfrac{1}{2}log(z+i) + \dfrac{1}{2}log(z-i) + logC"


b.) "\int \dfrac{2x+1}{x^2(x^2+1)} dx"


Performing partial fraction decomposition:


"\dfrac{2x+1}{x^2(x^2+1)} = \dfrac{A}{x} + \dfrac{B}{x^2}+ \dfrac{Cx+D}{x^2+1}"


"2x+1 = x^2(Cx+D) + x(x^2+1)A + (x^2+1)B"


"2x+1 = x^3(A+C) + x^2(B+D) + xA+B"


After equating coefficients we get,


"A= 2 , B=1 ,C = -2 \hspace{2mm}and \hspace{2mm} D = -1"


Hence,


"= \int (\dfrac{-2x-1}{x^2+1}+ \dfrac{2}{x}+ \dfrac{1}{x^2})dx"


Apply linearity:


"= -\int \dfrac{2x+1}{x^2+1}dx + 2\int \dfrac{1}{x}dx + \int \dfrac{1}{x^2}dx"


"= -\int \dfrac{2x}{x^2+1}dx - \int \dfrac{1}{x^2+1}dx + 2\int \dfrac{1}{x}dx + \int \dfrac{1}{x^2}dx"


"= -log(x^2+1) - tan^{-1}x + 2logx - \dfrac{1}{x}+ C"